OpenStudy (anonymous):

Perform the indicated operations: (7x^3-5x+2)-(5^3-4x^2+6x-7)

OpenStudy (hihi67):

@hartnn

OpenStudy (hihi67):

@madrockz

OpenStudy (anonymous):

We can + or - and *or divided the same sentence . Same sentence are should be have this : 1 . the same words for example x and x are same but x and y are not . 2. have same power for example : x^2 and x^2 . x^5 and x^6 are not same .or x^6 and y^6 are not . So the sames in this are : -6x and -5x caz in the both we have x . So can sum them . But first we have : (7x^3-5x+2)-(5^3-4x^2+6x-7)=(7x^3-5x+2)+(-5^3+4x^2-6x+7) K ? @BasedBoyIsaiah

OpenStudy (anonymous):

Okay I'm getting it

OpenStudy (anonymous):

So same them : (7x^3-5x+2)+(-5^3+4x^2-6x+7)=-5x-6x+(7x^3+2)+(-5^3+4x^2+7)= -11x+(7x^3+2)+(-5^3+4x^2+7) And for numbers : -11x+2+7+7x^3+-5^3+4x^2=-11x+9+7x^3+-5^3+4x^2=11x-125+9+7x^3 And the answer is : 11x-125+9+7x^3 = 11x-116+7x^3 Got it ? Are you sure it is -5^3? I think it s -5x^3 how ever there is no problem .

OpenStudy (anonymous):

@BasedBoyIsaiah :Got it ?